OpenStudy (anonymous):

The width of a rectangle is 12 feet shorter than twice its length. Find the width, and only the width, of the rectangle if the area is 224 square feet.

OpenStudy (anonymous):

Ok, so the formula I keep ending up with is \[w =\frac{ 1 }{ 2 }L +4\]

OpenStudy (anonymous):

but the equation \[134=L \left( \frac{ 1 }{ 2 }l +4 \right)\]

OpenStudy (anonymous):

\[134=\frac{ 1 }{ 2 }L ^{2}\]

hartnn (hartnn):

1/2 L^2 +4L

OpenStudy (anonymous):

134=(4+(w/2))w

OpenStudy (anonymous):

\[268=L ^{2}\]

hartnn (hartnn):

you did't distribute!

hartnn (hartnn):

\(134=L \left( \frac{ 1 }{ 2 }L +4 \right)= (1/2)L^2 +L\times 4= L^2/2+4L\)

hartnn (hartnn):

you know there's a simpler way ...

OpenStudy (anonymous):

let length =l and width = b \[b=\frac{1}{2}l+4\] Area of rectangle = lb \[134=l(\frac{1}{2}l+4)\] \[134=\frac{1}{2}l^2+4l \rightarrow 268 =l^2+8 l \rightarrow l^2+8 l -268=0\] Now factorize it and u will find the ans

hartnn (hartnn):

w= 1/2 L + 4 gives L = (2w-4) Area = Lw = (2w-4)w = 134 you directly get w from this

OpenStudy (anonymous):

yeah, I ended up getting \[0=L ^{2}+8L -268\] but I couldn't get that to factor

hartnn (hartnn):

***2 (w-4) Area = Lw = 2(w-4)w = 134

hartnn (hartnn):

w^2-4w-68 = 0 got this ? or you want to continue with your work of finding L first ??

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

Well now I got \[w ^{2}-4W -67=0\] so yes, i'm confused again

hartnn (hartnn):

67 ? or 68 ?

hartnn (hartnn):

that has irrational roots, not integer

hartnn (hartnn):

yeah, its 67, sorry

OpenStudy (anonymous):

We did several problems like this. It usually came out to a negative and positive integer. Since the length can't be negative, we just chose the positive one. So, I don't think I should have no solution for this, but i'm not sure. Math professors just love to give you wacky problems that they didn't go over.

hartnn (hartnn):

w^2-4w-67=0 has 2 solutions, one positive and one negative.....

hartnn (hartnn):

but both irrational

OpenStudy (anonymous):

so i'm wondering whether I should say no rational numbers or just state the formula w=1/2L + 4

OpenStudy (anonymous):

we don't deal with irrational or imaginary in this class.

OpenStudy (anonymous):

you lost me after \[w ^{2}-4w -67=0\]

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

heard of that quadratic formula ?

OpenStudy (anonymous):

I know from mentally doing this equation that the answer is somewhere between 11 and 13

OpenStudy (anonymous):

great!